OpenStudy (xapproachesinfinity):

hey need some company to do this problem find all continuous positive functions for 0<=x<=1, and \[\int_0^1 f(x)dx=1\] \[\int_0^1 xf(x)dx=\alpha\] \[\int_0^1 x^2f(x)dx=\alpha^2\] alpha is a real number

OpenStudy (xapproachesinfinity):

my first impression do this \[\int_0^1(x^2-x-1)f(x)=\alpha^2-\alpha-1\]

OpenStudy (anonymous):

i'm stumped my first idea was parts, but that doesn't seem to help

OpenStudy (xapproachesinfinity):

i thought parts didn't seem to lead me anywhere

OpenStudy (kainui):

That's clever, it looks like you're headed towards some kind of fibonacci thing @xapproachesinfinity I was also thinking about IBP. Have you tried taylor series?

OpenStudy (xapproachesinfinity):

i was looking into something of that nature @Kainui series didn't click to me

OpenStudy (kainui):

Maybe if playing with series falls through... LOL \[\int_0^1 x^2 f(x) dx = \int_0^1 \int_0^1 xyf(x)f(y)dxdy\]

OpenStudy (xapproachesinfinity):

gotta go for now, i will work on it more tomorrow

OpenStudy (kainui):

Yeah good luck, this is an interesting problem I'd like to finish it.